Fast Care is a health care franchise that functions as a primary family health clinic, seeing unscheduled patients 24 hours per day. A corporate office management engineering study has shown that the clinic was experiencing some dips in volume in the mid-afternoon hours. To increase volume, efficiney, and revenues, the clinic administrator contracted with the area high schools to provide after-school physicals for the sports teams. The initial agreement was that Fast Care would charge the market average which is $150 per visit. Fixed costs were $50,000 and variable costs were $55 per physical. Although this strategy proved somewhat successful, gross profit margin lagged behind the corporate expectations. To improve its margin, the clinic is considering increasing the visit price to $175. Fast Care projects that this price increase will cause the high schools to send their athletes to other providers which could cause a 33% decrease in volume. Last year, Fast Care performed 1,500 visits. If the program closes down entirely, all $50,000 in fixed costs would be saved.

a. Assuming that this price increase would decrease the number of patients by 33%, what should Fast Care do.

b What price would Fast Care have to charge to make up for the loss of patients

c. What if only 40% of the fixed costs are avoidable? What decision should Fast Care make in that case.
Answer #1
Solution:
a If Fast Care increases the price it will have the following 2 consequences
1 Patients visiting after 33% decrease = 1500*(1-0.33) = 1,005
Revenue as per new rate:
Particulars Amount
Collection (1,005*175)     175,875
Less: Variable Cost (1,005*55)       55,275
Less: Fixed Cost       50,000
Profit (Collection- Variable cost- Fixed Cost)       70,600
2 If the program closes down entirely then saving in cost = 50,000
So, Fast Care should go with option 1 as it is earning profit of 70,600 which is more than the saving in option 2
b If Fast Care would not have lost the patients then it's earnings would be
Particulars Amount
Collection (1,500*175)     262,500
Less: Variable Cost (1,500*55)       82,500
Less: Fixed Cost       50,000
Profit (Collection- Variable cost- Fixed Cost)     130,000
Let the new fee to be charged from each patient be 'x'
1005 (x-55) - 50,000 = 130,000
1005x - 55,275 - 50,000 = 130,000
1005x = 130,000+55,275+50,000
x = 235,275/1005
x = 235
If Fast Care wants to earn the same revenue as it was earning when rate was 150
Then, revenue at that rate was
Particulars Amount
Collection (1,500*150)     225,000
Less: Variable Cost (1,500*55)       82,500
Less: Fixed Cost       50,000
Profit (Collection- Variable cost- Fixed Cost)       92,500
Let the new fee to be charged from each patient be 'x'
1005 (x-55) - 50000 = 92,500
1005x - 55,275 - 50,000 = 92,500
1005x = 92,500+55,275+50,000
x = 197,775/1005
x = 197
c If only 40% of 50,000 i.e 20,000 is avoidable, then the company should not close it's program.